User name ulmer
Log entry time 16:01:34 on October22,2002
Entry number 88802
keyword=Day Shift Summary
08:00-08:25 Finish boiling studies with 15 cm LH2 (up to 110 uA).
This did not include varying the spot size. We should do a couple of spots at the highest current when we return to the f00 kinematics as part of the production running.
08:25-08:30 Harp scan 1884 - see Halog 88772.
08:30: Beam taken away for machine development.
10:20: Harp #6 may now be working: see Halog 88773
12:30: Bob and Bodo connected up unprescaled T1/T3/T5 to an 1877 TDC (see Halog 88777).
1345: Optics Runs 2675, 2676 with Optics Target
1440: Target change to LD2
Run 2677, LD2 target, 6 uA, 4x4 raster.
1528: Run 2678, LD2 target, 110 uA, 2.8x2.8 raster nominal (4x4 at tgt), 132k events
1543: Run 2679, LD2 target, 110 uA,